Determine the notation ("\mu,\sigma,\rho,\sigma^2" ), symbols (= , "\ne" , <, >, "\le, \ge" ), and value of the parameter:

  1. Average salary of University of Eastern Philippines (UEP) graduates is at most ₱ 324, 000.
  2. The standard deviation of students riding a tricycle from Catarman to Dalakit is 1.5
  3. The Department of Education offers a mean of 14 days of vacation for students from May 2-13.
  4. Passing rate of students in Statistics and Probability Philippines is below 60%.
  5. Mean number of teachers in Catarman National High School is at most 1.9


Notations Symbols Value

  1. "\le"
  2. =
  3. 14
  4. .60
  5. "\mu"
1
Expert's answer
2022-06-06T23:44:14-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & Notations & Symbols & Value \\ \hline
 1 & \mu & \le & 324000 \\
 \hdashline
 2 & \sigma & = & 1.5 \\
 \hdashline
 3 & \mu & = & 14 \\
 \hdashline
 4 & \rho & < & 0.6\\
 \hdashline
 5 & \mu & \le & 1.9\\
\end{array}"